Find Real GDP, assuming the GDP price index is 113.4 given:
Corporate Profits 31
Consumption of Fixed Capital 35
Gross Private Domestic Investment 106
Personal Taxes 56
Personal Saving 82
Government Purchases 128
Imports 85
Interest 59
Net Foreign Factor Income Earned 2
Undistributed Corporate Profits 5
Exports 67
Personal Consumption Expenditures 377
Statistical Discrepancy 10
Taxes on Production and Imports 25
Social Security Contributions 39
Transfer Payments 48
Rents 13
Proprietors' Income 33
Corporate Income Taxes 6
Compensation of Employees 389
1
Expert's answer
2012-11-09T05:38:41-0500
GDP = Personal Consumption Expenditures + Gross Private Domestic Investment + Government Purchases +
(Exports - Imports) = 377 + 106 + 128 + (67 - 85) = 593
Real GDP = 593/1.134 = $522.9